# Improving Stance Detection by Leveraging Measurement Knowledge from Social Sciences: A Case Study of Dutch Political Tweets and Traditional Gender Role Division

### v2

@@ 4 Conclusion & Discussion
In this paper, we apply SD to a dataset of tweets from official party accounts in the Netherlands between 2017 and 2021, with a focus on detecting stances towards traditional gender role division. We also investigate whether using established survey instruments can be helpful for SD.
In Experiment 1, we are faced the challenge of lacking labelled SD datasets in Dutch. We thus opt to reformulate our SD task as TER, where we train a BERTje model on a Dutch TER dataset and subsequently apply it to zero-shot SD. We find using survey items consistently outperforms using a simple hypothesis for tweet-level SD. However, our model fails to achieve good SD on the party level. In Experiment 2, we further test the benefit of using survey instruments by repeating the tweet-level analysis from Experiment 1 using prompts with OpenAI’s GPT-3.5 Turbo API. We again find using survey questions beneficial for SD of traditional gender role division.
Our study has limitations. For instance, we focus on a single SD target and only a subset of the political tweet dataset. Nevertheless, we hope that our study will inspire SD researchers to explore the use of survey instruments for SD research further. For future research, we suggest, among others, converting existing SD datasets into TER-compatible formats for model training, developing more general TER datasets (beyond concrete scenes), using multilingual models (where datasets in different languages may be combined), and comparing different strategies for constructing TER hypotheses from survey items.
A Additional Reproducibility Information
@@
For the BERTje model, we use the same hyperparameters and search strategy as Wijnholds and Moortgat (2021), including:
@@
B The Survey Instruments and Hypotheses
@@
C Examples of Tweets of Different Stances towards Traditional Gender Role Division

### v3

@@ 4 Conclusion & Discussion
In this paper, we apply stance detection to a dataset of tweets from official party accounts in the Netherlands between 2017 and 2021, with a focus on detecting stances towards traditional gender role division. We also investigate whether using established survey instruments can be helpful for stance detection.
In Experiment 1, we are faced with the challenge of lacking labelled stance detection datasets in Dutch. We thus opt to reformulate our stance detection task as TER, where we train a BERTje model on a Dutch TER dataset and subsequently apply it to zero-shot stance detection. We find using survey items consistently outperforms using a simple hypothesis for tweet-level stance detection. However, our model fails to achieve good stance detection on the party level. In Experiment 2, we further test the benefit of using survey instruments by repeating the tweet-level analysis from Experiment 1 using prompts with OpenAI’s GPT-3.5 Turbo API. We again find using survey questions beneficial for stance detection of traditional gender role division.
Our study has limitations. For instance, we focus on a single stance detection target and only a subset of the political tweet dataset. Including more stance detection targets and more datasets is necessary to further support the generalisability of our findings. Nevertheless, we hope that our study will inspire stance detection researchers to explore the use of survey instruments for stance detection research further. Furthermore, our second experiment, which utilised GPT-3.5, is now somewhat outdated given the recent emergence of more capable large language models like GPT-4 and GPT-4o. We also do not incorporate with other prompting engineering techniques such as chain-of-thought and few-short prompting, which have the potential to improve the model’s task performance.
For future research, we suggest, among others, converting existing stance detection datasets into TER-compatible formats for model training, developing more general TER datasets (beyond concrete scenes), using multilingual models (where datasets in different languages may be combined), and comparing different strategies for constructing TER hypotheses from survey items.
{credits}
4.0.1 Acknowledgements
We thank Dong Nguyen for her constructive and helpful feedback. This work was partially supported by the Dutch Research Council (NWO): grant number: VI.Vidi.195.152 to D. L. Oberski.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Additional Reproducibility Information
@@
For the BERTje model, we use the same hyperparameters and search strategy as [24], including:
@@
Appendix 0.B The Survey Instruments and Hypotheses
@@
Appendix 0.C Examples of Tweets of Different Stances towards Traditional Gender Role Division